# Convergence and long-time behavior of finite volumes for a generalized Poisson-Nernst-Planck system with cross-diffusion and size exclusion

### v1

@@ 4.2. Identification of the limit
The regularity requirements on $U$ and $\phi$ have already been checked in Proposition 4.1. Therefore it only remains to verify that the weak formulations (16) and (17) hold true. The case of Poisson equation is classical. It will not be detailed here, and we refer to [14, Proposition 4.2] for a synthetic proof. We rather focus our attention on the derivation of (17).
Let $\varphi\in C^{\infty}_{c}(\mathbb{R}_{+}\times\overline{\Omega})$, then, for some admissible mesh $(\mathcal{T}_{\ell},\mathcal{E}_{\ell},(x_{K})_{K\in\mathcal{T}_{\ell}})$ (we remove the subscript $\ell$ when possible for legibility), denote by $\varphi_{K}^{n}=\varphi(t^{n},x_{K})$. Multiplying (25) by $\tau^{n}\varphi_{K}^{n-1}$ and summing over $n\geq 1$ and $K\in\mathcal{T}$ yields, as for (56):
(68) $\sum_{n=1}^{N}\sum_{K\in\mathcal{T}}m_{K}(u_{i,K}^{n}-u_{i,K}^{n-1})\varphi_{K}^{n-1}=\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\text{int}}}F_{i,K\sigma}^{n}(\varphi_{K\sigma}^{n-1}-\varphi_{K}^{n-1}).$
Since $\varphi$ is compactly supported in time, $\varphi_{K}^{n}=0$ for $n$ large enough. Therefore, the left-hand side of (68) rewrites
@@
converges in $L^{1}(\mathbb{R}_{+}\times\Omega)$ towards $\partial_{t}\varphi$. Therefore, using (60), we get that
(69) $\sum_{n=1}^{N}\tau^{n}\sum_{K\in\mathcal{T}}m_{K}u_{i,K}^{n}\frac{\varphi_{K}^{n-1}-\varphi_{K}^{n}}{\tau^{n}}\underset{\ell\to+\infty}{\longrightarrow}-\iint_{\mathbb{R}_{+}\times\Omega}u_{i}\partial_{t}\varphi.$
@@
(70) $-\sum_{K\in\mathcal{T}}m_{K}u_{i,K}^{0}\varphi_{K}^{0}\underset{\ell\to+\infty}{\longrightarrow}-\int_{\Omega}u_{i}^{0}\varphi(0,\cdot),$
@@
Following [21] (see [19] for a practical example), one can reconstruct a second approximate gradient operator ^$\widehat{\nabla}_{\mathcal{T}}$ mapping the set of piecewise constant functions in space and time to $\mathbb{R}^{d}$ such that
^$\widehat{\nabla}_{\mathcal{T}}\tilde{\varphi}_{\mathcal{T},\tau}(t,x)\cdot n_{K\sigma}=\frac{\varphi_{L}^{n}-\varphi_{K}^{n}}{d_{\sigma}}\quad\text{if}\;(t,x)\in[t^{n-1},t^{n})\times\omega_{\sigma},\quad\sigma=K|L\in\mathcal{E}_{\text{int}},$
@@
^$\widehat{\nabla}_{\mathcal{T}}\tilde{\varphi}_{\mathcal{T},\tau}\underset{\ell\to+\infty}{\longrightarrow}\nabla\varphi\quad\text{strongly in }L^{2}(\mathbb{R}_{+}\times\Omega)^{d}.$
The right-had side of (68) then simply boils down to
$\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\text{int}}}F_{i,K\sigma}^{n}(\varphi_{K\sigma}^{n-1}-\varphi_{K}^{n-1})=\iint_{\mathbb{R}_{+}\times\Omega}F_{i,\mathcal{E},\tau}\cdot\widehat{\nabla}_{\mathcal{T}}\tilde{\varphi}_{\mathcal{T},\tau}.$
We have shown in Lemma 3.8 that $\left(F_{i,\mathcal{E}_{\ell},\tau_{\ell}}\right)_{\ell\geq 1}$ is uniformly bounded in $L^{2}_{\text{loc}}(\mathbb{R}_{+}\times\overline{\Omega})^{d}$. Therefore, it converges (up to a subsequence) towards some vector field $F_{i}$ weakly in $L^{2}_{\text{loc}}(\mathbb{R}_{+}\times\overline{\Omega})$. We can then pass to the limit in the right-hand side of (68) to obtain
(71) $\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\text{int}}}F_{i,K\sigma}^{n}(\varphi_{K\sigma}^{n-1}-\varphi_{K}^{n-1})\underset{\ell\to+\infty}{\longrightarrow}\iint_{\mathbb{R}_{+}\times\Omega}F_{i}\cdot\nabla\varphi.$
@@
where $F_{i,K\sigma}^{\text{conv},n}$ and $F_{i,K\sigma}^{\text{diff},n}$ are as in the proof of Lemma 3.8. As both $F_{i,\mathcal{E},\tau}^{\text{conv}}$ and $F_{i,\mathcal{E},\tau}^{\text{diff}}$ have been shown to be bounded in $L^{2}_{\text{loc}}(\mathbb{R}_{+}\times\overline{\Omega})^{d}$, one has, up to a subsequence,
(72) $F_{i,\mathcal{E},\tau}^{\text{diff}}\underset{\ell\to+\infty}{\longrightarrow}F_{i}^{\text{diff}}\quad\text{and}\quad F_{i,\mathcal{E},\tau}^{\text{conv}}\underset{\ell\to+\infty}{\longrightarrow}F_{i}^{\text{conv}}\quad\text{weaklyin }L^{2}_{\text{loc}}(\mathbb{R}_{+}\times\overline{\Omega})^{d}.$
@@
$\eta_{i,\sigma}^{n}=\begin{cases}\frac{u_{i,K}^{n}u_{0,L}^{n}+u_{i,L}^{n}u_{0,K}^{n}}{2}&\text{if}\;\sigma=K|L\in\mathcal{E}_{\text{int}},\\ u_{i,K}^{n}u_{0,K}^{n}&\text{if}\;\sigma\in\mathcal{E}_{K}\cap\mathcal{E}_{\text{ext}},\end{cases}$
@@
$\tilde{\eta}_{i,\sigma}^{n}=\begin{cases}\frac{u_{i,K}^{n}u_{0,K}^{n}+u_{i,L}^{n}u_{0,L}^{n}}{2}&\text{if}\;\sigma=K|L\in\mathcal{E}_{\text{int}},\\ u_{i,K}^{n}u_{0,K}^{n}&\text{if}\;\sigma\in\mathcal{E}_{K}\cap\mathcal{E}_{\text{ext}},\end{cases}$
and then $\eta_{i,\mathcal{E},\tau}(t,x)=\eta_{i,\sigma}^{n}$ and $\tilde{\eta}_{i,\mathcal{E},\tau}(t,x)=\eta_{i,\sigma}^{n}$ if $(t,x)\in(t^{n-1},t^{n}]\times\omega_{\sigma}$. In view of (53) and of the definition (59) of the inflated gradient, the convective part of the inflated approximate flux rewrites
@@
(74) $\eta_{i,\mathcal{E},\tau}-\tilde{\eta}_{i,\mathcal{E},\tau}\underset{\ell\to+\infty}{\longrightarrow}0\quad\text{a.e. in }\mathbb{R}_{+}\times\Omega$
@@
(75) $\tilde{\eta}_{i,\mathcal{E},\tau}-u_{0,\mathcal{T},\tau}u_{i,\mathcal{T},\tau}\underset{\ell\to+\infty}{\longrightarrow}0\quad\text{a.e. in }\mathbb{R}_{+}\times\Omega$
@@
$\eta_{i,\mathcal{E},\tau}\underset{\ell\to+\infty}{\longrightarrow}u_{0}u_{i}\quad\text{a.e. in }\mathbb{R}_{+}\times\Omega,$
and thus strongly in $L^{2}_{\text{loc}}(\mathbb{R}_{+}\times\overline{\Omega})$ thanks to the uniform bound $0\leq u_{i,\mathcal{T},\tau}\leq 1$. Bearing in mind (67), this allows to pass to the weak limit in (73) and to recover that
@@
$\left\|\eta_{i,\mathcal{E},\tau}-\tilde{\eta}_{i,\mathcal{E},\tau}\right\|_{L^{2}((0,t^{N})\times\Omega)}^{2}=\frac{1}{4}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\text{int}}}m_{\omega_{\sigma}}(u_{i,K}^{n}-u_{i,L}^{n})^{2}(u_{0,K}^{n}-u_{0,L}^{n})^{2}.$
@@
$\left\|\eta_{i,\mathcal{E},\tau}-\tilde{\eta}_{i,\mathcal{E},\tau}\right\|_{L^{2}((0,t^{N})\times\Omega)}^{2}\leq\frac{h_{\mathcal{T}}^{2}}{d}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\text{int}}}a_{\sigma}(u_{0,K}^{n}-u_{0,L}^{n})^{2}.$
We can make use of estimate (47) to get that $\eta_{i,\mathcal{E},\tau}-\tilde{\eta}_{i,\mathcal{E},\tau}$ tends to 00 in $L^{2}_{\text{loc}}(\mathbb{R}_{+}\times\overline{\Omega})$, hence almost everywhere up to the extraction of yet another subsequence, whence (74).
@@
$\left\|\tilde{\eta}_{i,\mathcal{E},\tau}-u_{0,\mathcal{T},\tau}u_{i,\mathcal{T},\tau}\right\|_{L^{2}((0,t^{N})\times\Omega)}^{2}{\leq}\frac{h_{\mathcal{T}}^{2}}{d}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\text{int}}}a_{\sigma}\left(u_{0,K}^{n}u_{i,K}^{n}-u_{0,L}^{n}u_{i,L}^{n}\right)^{2}$
which tends to 00 thanks to (49).
It finally remains to identify the limit of $F_{i,\mathcal{E},\tau}^{\text{diff}}$ as $F_{i}^{\text{diff}}$. To this end, remark first that, for $\sigma=K|L\in\mathcal{E}_{\text{int}}$ and $n\geq 1$, one has
@@
$\gamma_{i,\mathcal{E},\tau}(t,x)=\begin{cases}\frac{1}{2}\left(u_{i,K}^{n}\sqrt{u_{0,K}^{n}}+u_{i,L}^{n}\sqrt{u_{0,L}^{n}}\right)&\text{if}\;(t,x)\in(t^{n-1},t^{n}]\times\omega_{\sigma},\;\sigma=K|L\in\mathcal{E}_{\text{int}},\\ u_{i,K}^{n}\sqrt{u_{0,K}^{n}}&\text{if}\;(t,x)\in(t^{n-1},t^{n}]\times\omega_{\sigma},\;\sigma\in\mathcal{E}_{K}\cap\mathcal{E}_{\text{ext}},\end{cases}$
and where, for $(t,x)\in(t^{n-1},t^{n}]\times\omega_{\sigma}$, $\sigma=K|L\in\mathcal{E}_{\text{int}},$
@@
Both $\mathcal{R}_{i,\mathcal{E},\tau}$ and $\mathcal{S}_{i,\mathcal{E},\tau}$ are assumed to vanish on +×$\mathbb{R}_{+}\times\omega_{\sigma}$ for all $\sigma\in\mathcal{E}_{\text{ext}}$. Then thanks to Proposition 3.3, one has
$\|\mathcal{S}_{i,\mathcal{E},\tau}\|_{L^{1}((0,t^{N})\times\Omega)^{d}}\leq C\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\text{int}}}m_{\sigma}\left(\phi_{K}^{n}-\phi_{L}^{n}\right)^{2}\leq Ch_{\mathcal{T}}t^{N}.$
@@
$\|\mathcal{R}_{i,\mathcal{E},\tau}\|_{L^{1}((0,t^{N})\times\Omega)^{d}}\leq D_{i}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\text{int}}}m_{\sigma}\left(\sqrt{u_{0,K}^{n}}-\sqrt{u_{0,L}^{n}}\right)^{2}\leq Ch_{\mathcal{T}}(1+t^{N}).$
The two last terms in (77) then tend to 00, while the first term tends to $-D_{i}\nabla[u_{0}u_{i}]$ thanks to (66). In view of the weak $L^{2}_{\text{loc}}(L^{2})$ convergence of $\nabla_{\mathcal{T}}\sqrt{u_{0,\mathcal{T},\tau}}$ towards $\nabla\sqrt{u_{0}}$, cf. (65), then it suffices to show that
$\gamma_{i,\mathcal{E},\tau}\underset{\ell\to+\infty}{\longrightarrow}u_{i}\sqrt{u_{0}}\quad\text{strongly in\; }L^{2}_{\text{loc}}(\mathbb{R}_{+}\times\overline{\Omega})$
@@
$\|\gamma_{i,\mathcal{E},\tau}-u_{i,\mathcal{T},\tau}\sqrt{u_{0,\mathcal{T},\tau}}\|_{L^{2}((0,t^{N})\times\Omega)}^{2}=\;$ $\frac{1}{2}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\text{int}}}m_{\omega_{\sigma}}\left(u_{i,K}^{n}\sqrt{u_{0,K}^{n}}-u_{i,L}^{n}\sqrt{u_{0,L}^{n}}\right)^{2}$

### v2

@@ 4.2. Identification of the limit
The regularity requirements on $U$ and $\phi$ have already been checked in Proposition 4.1. Therefore it only remains to verify that the weak formulations (16) and (17) hold true. The case of the Poisson equation is classical. It will not be detailed here, and we refer to [14, Proposition 4.2] for a synthetic proof. We rather focus our attention on the derivation of (17).
Let $\varphi\in C^{\infty}_{c}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\overline{\Omega})$, then, for some admissible mesh $(\mathcal{T}_{\ell},\mathcal{E}_{\ell},(x_{K})_{K\in\mathcal{T}_{\ell}})$ (we remove the subscript $\ell$ when possible for legibility), denote by $\varphi_{K}^{n}=\varphi(t^{n},x_{K})$. Multiplying (25) by $\tau^{n}\varphi_{K}^{n-1}$ and summing over $n\geq 1$ and $K\in\mathcal{T}$ yields, as for (56):
(68) $\sum_{n=1}^{N}\sum_{K\in\mathcal{T}}m_{K}(u_{i,K}^{n}-u_{i,K}^{n-1})\varphi_{K}^{n-1}=\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}F_{i,K\sigma}^{n}(\varphi_{K\sigma}^{n-1}-\varphi_{K}^{n-1}).$
Since $\varphi$ is compactly supported in time, $\varphi_{K}^{n}=0$ for $n$ large enough. Therefore, the left-hand side of (68) is rewritten as
@@
converges in $L^{1}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega)$ towards $\partial_{t}\varphi$. Therefore, using (60), we get that
(69) $\sum_{n=1}^{N}\tau^{n}\sum_{K\in\mathcal{T}}m_{K}u_{i,K}^{n}\frac{\varphi_{K}^{n-1}-\varphi_{K}^{n}}{\tau^{n}}\underset{\ell\to+\infty}{\longrightarrow}-\iint_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega}u_{i}\partial_{t}\varphi.$
@@
(70) $\sum_{K\in\mathcal{T}}m_{K}u_{i,K}^{0}\varphi_{K}^{0}\underset{\ell\to+\infty}{\longrightarrow}-\int_{\Omega}u_{i}^{0}\varphi(0,\cdot),$
@@
Following [23] (see [20] for a practical example), one can reconstruct a second approximate gradient operator $\widehat{\nabla}_{\mathcal{T}}$ mapping the set of piecewise constant functions in space and time to $\mathbb{R}^{d}$ such that
$\widehat{\nabla}_{\mathcal{T}}\tilde{\varphi}_{\mathcal{T},\tau}(t,x)\cdot n_{K\sigma}=\frac{\varphi_{L}^{n}-\varphi_{K}^{n}}{d_{\sigma}}\quad\text{if}\;(t,x)\in[t^{n-1},t^{n})\times\omega_{\sigma},\quad\sigma=K|L\in\mathcal{E}_{\rm int},$
@@
$\widehat{\nabla}_{\mathcal{T}}\tilde{\varphi}_{\mathcal{T},\tau}\underset{\ell\to+\infty}{\longrightarrow}\nabla\varphi\quad\text{strongly in }L^{2}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega)^{d}.$
The right-hand side of (68) then simply boils down to
$\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}F_{i,K\sigma}^{n}(\varphi_{K\sigma}^{n-1}-\varphi_{K}^{n-1})=\iint_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega}F_{i,\mathcal{E},\tau}\cdot\widehat{\nabla}_{\mathcal{T}}\tilde{\varphi}_{\mathcal{T},\tau}.$
We have shown in Lemma 3.9 that $\left(F_{i,\mathcal{E}_{\ell},\tau_{\ell}}\right)_{\ell\geq 1}$ is uniformly bounded in $L^{2}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\overline{\Omega})^{d}$. Therefore, it converges (up to a subsequence) towards some vector field $F_{i}$ weakly in $L^{2}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\overline{\Omega})$. We can then pass to the limit in the right-hand side of (68) to obtain
(71) $\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}F_{i,K\sigma}^{n}(\varphi_{K\sigma}^{n-1}-\varphi_{K}^{n-1})\underset{\ell\to+\infty}{\longrightarrow}\iint_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega}F_{i}\cdot\nabla\varphi.$
@@
where $F_{i,K\sigma}^{\text{conv},n}$ and $F_{i,K\sigma}^{\text{diff},n}$ are as in the proof of Lemma 3.9. As both $F_{i,\mathcal{E},\tau}^{\text{conv}}$ and $F_{i,\mathcal{E},\tau}^{\text{diff}}$ have been shown to be bounded in $L^{2}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\overline{\Omega})^{d}$, one has, up to a subsequence,
(72) $F_{i,\mathcal{E},\tau}^{\text{diff}}\underset{\ell\to+\infty}{\longrightarrow}F_{i}^{\text{diff}}\quad\text{and}\quad F_{i,\mathcal{E},\tau}^{\text{conv}}\underset{\ell\to+\infty}{\longrightarrow}F_{i}^{\text{conv}}\quad\text{weakly in }L^{2}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\overline{\Omega})^{d}.$
@@
$\eta_{i,\sigma}^{n}=\begin{cases}\frac{u_{i,K}^{n}u_{0,L}^{n}+u_{i,L}^{n}u_{0,K}^{n}}{2}&\text{if}\;\sigma=K|L\in\mathcal{E}_{\rm int},\\ u_{i,K}^{n}u_{0,K}^{n}&\text{if}\;\sigma\in\mathcal{E}_{K}\cap\mathcal{E}_{\text{ext}},\end{cases}$
@@
$\tilde{\eta}_{i,\sigma}^{n}=\begin{cases}\frac{u_{i,K}^{n}u_{0,K}^{n}+u_{i,L}^{n}u_{0,L}^{n}}{2}&\text{if}\;\sigma=K|L\in\mathcal{E}_{\rm int},\\ u_{i,K}^{n}u_{0,K}^{n}&\text{if}\;\sigma\in\mathcal{E}_{K}\cap\mathcal{E}_{\text{ext}},\end{cases}$
and then $\eta_{i,\mathcal{E},\tau}(t,x)=\eta_{i,\sigma}^{n}$ and $\tilde{\eta}_{i,\mathcal{E},\tau}(t,x)=\eta_{i,\sigma}^{n}$ if $(t,x)\in(t^{n-1},t^{n}]\times\omega_{\sigma}$. In view of (53) and of the definition (59) of the inflated gradient, the convective part of the inflated approximate flux is rewritten as
@@
(74) $\eta_{i,\mathcal{E},\tau}-\tilde{\eta}_{i,\mathcal{E},\tau}\underset{\ell\to+\infty}{\longrightarrow}0\quad\text{a.e. in }{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega$
@@
(75) $\tilde{\eta}_{i,\mathcal{E},\tau}-u_{0,\mathcal{T},\tau}u_{i,\mathcal{T},\tau}\underset{\ell\to+\infty}{\longrightarrow}0\quad\text{a.e. in }{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega$
@@
$\eta_{i,\mathcal{E},\tau}\underset{\ell\to+\infty}{\longrightarrow}u_{0}u_{i}\quad\text{a.e. in }{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega,$
and thus strongly in $L^{2}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\overline{\Omega})$ thanks to the uniform bound $0\leq u_{i,\mathcal{T},\tau}\leq 1$. Bearing in mind (67), this allows to pass to the weak limit in (73) and to recover that
@@
$\left\|\eta_{i,\mathcal{E},\tau}-\tilde{\eta}_{i,\mathcal{E},\tau}\right\|_{L^{2}((0,t^{N})\times\Omega)}^{2}=\frac{1}{4}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}m_{\omega_{\sigma}}(u_{i,K}^{n}-u_{i,L}^{n})^{2}(u_{0,K}^{n}-u_{0,L}^{n})^{2}.$
@@
$\left\|\eta_{i,\mathcal{E},\tau}-\tilde{\eta}_{i,\mathcal{E},\tau}\right\|_{L^{2}((0,t^{N})\times\Omega)}^{2}\leq\frac{h_{\mathcal{T}}^{2}}{d}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}(u_{0,K}^{n}-u_{0,L}^{n})^{2}.$
We can make use of estimate (47) to get that $\eta_{i,\mathcal{E},\tau}-\tilde{\eta}_{i,\mathcal{E},\tau}$ tends to 0 in $L^{2}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\overline{\Omega})$, hence almost everywhere up to the extraction of yet another subsequence, whence (74).
@@
$\left\|\tilde{\eta}_{i,\mathcal{E},\tau}-u_{0,\mathcal{T},\tau}u_{i,\mathcal{T},\tau}\right\|_{L^{2}((0,t^{N})\times\Omega)}^{2}{\leq}\frac{h_{\mathcal{T}}^{2}}{d}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}\left(u_{0,K}^{n}u_{i,K}^{n}-u_{0,L}^{n}u_{i,L}^{n}\right)^{2}$
which tends to 0 thanks to (49).
It finally remains to identify the limit of $F_{i,\mathcal{E},\tau}^{\text{diff}}$ as $F_{i}^{\text{diff}}$. To this end, remark first that, for $\sigma=K|L\in\mathcal{E}_{\rm int}$ and $n\geq 1$, one has
@@
$\gamma_{i,\mathcal{E},\tau}(t,x)=\begin{cases}\frac{1}{2}\left(u_{i,K}^{n}\sqrt{u_{0,K}^{n}}+u_{i,L}^{n}\sqrt{u_{0,L}^{n}}\right)&\text{if}\;(t,x)\in(t^{n-1},t^{n}]\times\omega_{\sigma},\;\sigma=K|L\in\mathcal{E}_{\rm int},\\ u_{i,K}^{n}\sqrt{u_{0,K}^{n}}&\text{if}\;(t,x)\in(t^{n-1},t^{n}]\times\omega_{\sigma},\;\sigma\in\mathcal{E}_{K}\cap\mathcal{E}_{\text{ext}},\end{cases}$
and where, for $(t,x)\in(t^{n-1},t^{n}]\times\omega_{\sigma}$, $\sigma=K|L\in\mathcal{E}_{\rm int},$
@@
Both $\mathcal{R}_{i,\mathcal{E},\tau}$ and $\mathcal{S}_{i,\mathcal{E},\tau}$ are assumed to vanish on ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\omega_{\sigma}$ for all $\sigma\in\mathcal{E}_{\text{ext}}$. Then thanks to Proposition 3.3, one has
$\|\mathcal{S}_{i,\mathcal{E},\tau}\|_{L^{1}((0,t^{N})\times\Omega)^{d}}\leq C\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}m_{\sigma}\left(\phi_{K}^{n}-\phi_{L}^{n}\right)^{2}\leq Ch_{\mathcal{T}}t^{N}.$
@@
$\|\mathcal{R}_{i,\mathcal{E},\tau}\|_{L^{1}((0,t^{N})\times\Omega)^{d}}\leq D_{i}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}m_{\sigma}\left(\sqrt{u_{0,K}^{n}}-\sqrt{u_{0,L}^{n}}\right)^{2}\leq Ch_{\mathcal{T}}(1+t^{N}).$
The two last terms in (77) then tend to 0, while the first term tends to $D_{i}\nabla[u_{0}u_{i}]$ thanks to (66). In view of the weak $L^{2}_{\rm loc}(L^{2})$ convergence of $\nabla_{\mathcal{T}}\sqrt{u_{0,\mathcal{T},\tau}}$ towards $\nabla\sqrt{u_{0}}$, cf. (65), then it suffices to show that
$\gamma_{i,\mathcal{E},\tau}\underset{\ell\to+\infty}{\longrightarrow}u_{i}\sqrt{u_{0}}\quad\text{strongly in\; }L^{2}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\overline{\Omega})$
@@
$\|\gamma_{i,\mathcal{E},\tau}-u_{i,\mathcal{T},\tau}\sqrt{u_{0,\mathcal{T},\tau}}\|_{L^{2}((0,t^{N})\times\Omega)}^{2}=\;$ $\frac{1}{2}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}m_{\omega_{\sigma}}\left(u_{i,K}^{n}\sqrt{u_{0,K}^{n}}-u_{i,L}^{n}\sqrt{u_{0,L}^{n}}\right)^{2}$
$\leq\;$ $\frac{h_{\mathcal{T}}^{2}}{d}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}\left(u_{i,K}^{n}\sqrt{u_{0,K}^{n}}-u_{i,L}^{n}\sqrt{u_{0,L}^{n}}\right)^{2}$